OpenStudy (anonymous):

Identify the solutions of 3x2-x-4=0

OpenStudy (anonymous):

The solutions would -1 and 4/3

OpenStudy (anonymous):

3x^2 - x - 4 = 0 (3x-4)(x+1) = 0 x = -1, 4/3

OpenStudy (anonymous):

Multiply 3 and the -4 which equals -12. Then find factors of -12 that will add together to make -x then split the -x up. Rewrite the equation ---> 3x² +3x - 4x -4 = 0 Then group terms like this (3x² +3x) - (4x -4) = 0 3x(x+1) - 4(x-1) 3x and -4 are now both being multiplied by x-1, so you can isolate that as a common factor (x+1)(3x-4) = 0 Either x+1 = 0 or 3x-4 = 0 So your solutions are x= (-1) and x=4/3